OpenStudy (anonymous):

Given f(x)=2x^2-1 and g(x)=7-1/2x^2, find the following expressions. a.(f•g) (4) b.(g•f) (2) c.(f•f) (1) d.(g•g) (0)

OpenStudy (anonymous):

do the dots mean composite or multiplication? it's hard to tell

OpenStudy (anonymous):

multiplication

OpenStudy (ahsome):

I'll help you do (A) \[f(x)=2x^2-1\]\[g(x)=7-\frac{1}{2x^2}\] Now, you want to find:\[4(f*g)\] Simply sub the values from before into it \[4((2x^2-1)(7-\frac{1}{2x^2}))\] \[\frac{4}{1}((\frac{2x^2-1}{1})(\frac{7}{1}-\frac{1}{2x^2}))\] Multiply the section with \(2x^2\) \[\frac{4}{1}*\frac{2x^2}{2x^2}((\frac{2x^2-1}{1}*\frac{2x^2}{2x^2})(\frac{7}{1}*\frac{2x^2}{2x^2}-\frac{1}{2x^2}))\] \[\frac{8x^2}{2x^2}((\frac{4x^4}{2x^2})(\frac{7x^2}{2x^2}-\frac{1}{2x^2}))\] \[8x^2((4x^4)(7x^2-1))\] You can go from there